OpenStudy (anonymous):

TRIG ID's cot x =1/tan OR cot x = cos x/sin x

OpenStudy (anonymous):

both?

OpenStudy (anonymous):

Both

OpenStudy (anonymous):

\[\frac{ 1+\cos(y) }{ 1+\sec(y) }\]

OpenStudy (anonymous):

simplified equals 1 + cos^2 (y) = sin^2 (y) right?

OpenStudy (anonymous):

No, if it was \[1-\cos ^{2}y\] then it would be\[\sin ^{2}y\] but I think it is just \[\cos ^{2}y\]. Not 100% sure though

OpenStudy (anonymous):

1 + cos^2 (y) = ??

OpenStudy (anonymous):

It doesn't simplify beyond that, at least to my knowledge.

OpenStudy (anonymous):

ok thanks for your time sir

OpenStudy (anonymous):

No problem!